OpenStudy (anonymous):

Which ordered pair is on the graph of the equation 2x + 5y = 4? A. (7, –2) B. (0, 1) C. (3, –2) D. (0, –1) I believe the answer is C. But I am behind in math honestly and can't afford to be wrong

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

I have 3 question left. Ill medal you for every question I ask. For your trouble.

OpenStudy (welshfella):

lets check out C here x = 3 and y = -2 2x + 5y = 4 so if LHS = 4 we know its C 2(3) + 5(-2) = 6 - 10 = -4 no its not C

OpenStudy (welshfella):

try A

OpenStudy (welshfella):

2(7) + 5(-2)

OpenStudy (anonymous):

A = 4 does it not?

OpenStudy (welshfella):

yes A is the one

OpenStudy (anonymous):

I hate math :/

OpenStudy (anonymous):

75%. Which is good for me, but considering not mastering in my school

OpenStudy (welshfella):

yes - you need to like maths if you want to major in it

OpenStudy (anonymous):

I want to be a therapist and get my Ph.D. I do not think Math will be used very much

OpenStudy (welshfella):

right

OpenStudy (anonymous):

Anyways! Enough about the American school system and how poorly it is handled. Thank you for your help!

OpenStudy (welshfella):

lol yw